OpenStudy (jane11509):

Sharon is the youngest in her family. Her sister Anne is 3 years older than she is. Her brother Tom is 5 years older than Anne. The sum of their ages is 62. How old is Sharon?

OpenStudy (anonymous):

Solve the following for s, a and t:\[\{s=a-3,s=t-5,a+s+t=62\} \]

OpenStudy (mathmath333):

\(\large\ sharon =x\) \(\large\ anne =x+3\) \(\large\ tom =5+(x+3)\) \(\large\ x+(x+3)+(5+x+3)=62\)

OpenStudy (anonymous):

If one has access to Mathematica, a press of the key pad Enter key will solve for s directly. Eliminate[ {s=a-3,s=t-5,s+a+t=62}, {a,t} ] s = 18

OpenStudy (jane11509):

thank you

OpenStudy (anonymous):

You are welcome and thank you for the medal.

OpenStudy (anonymous):

i got it wrong and it said it is 17!

OpenStudy (anonymous):

Sorry about that. A new formulation is shown below: Solve[{s = a - 3, t = a + 5, a + s + t = 62}, {s, a, t}] {s -> 17, a -> 20, t -> 25}

OpenStudy (anonymous):

ok thanks!